Each of the 4 walls in the dining room measures 12 5/8 feet. How much chair rail must be purchased to install the chair rail on all four walls. Disregard any openings.

(Express your answer in feet using a mixed number. Use only one space between the whole number and the fraction.)

1
Expert's answer
2020-12-20T15:40:52-0500


total Length=101/8 x4=101/2 feet